The following table gives the number of pints of type A blood used at Damascus Hospital in the past 6 weeks 

Week OfPints Used
August 31360
September 7370
September 14410
September 21381
September 28366
October 5378

 a) The forecasted demand for the week of October 12 using a 3-week moving average = _______  pints .

 b) Using a 3-week weighted moving average, with weights of 0.15, 0.30, and 0.55, using 0.55 for the most recent week, the forecasted demand for the week of October 12 = _______ points.  (round your response to two decimal places and remember to use the weights in appropriate order the largest weight applies to most recent period and smallest weight applies to oldest period.) 

 c) lf the forecasted demand for the week of August 31 is 360 and α = 0.20, using exponential smoothing develop the forecast for each of the weeks with the known demand and the forecast for the week of October 12 .

Week OfPints UsedForecast for this Date
August 31360360
September 7370360.00
September 14410362.00
September 21381
September 28366
October 5378
October 12-
Answer #1

a) Using a 3-week moving average method, the forecast would be

F(October12) = {Actual(October5) + Actual(September28) + Actual(September21)}/3

= (378 + 366 + 381)/3

= 375

b) Using a 3-week weighted moving average method, the forecast would be:

F(October 12) = 0.55*Actual(October 5) + 0.30*Actual(September 28) + 0.15*Actual(September 21)

= 0.55*378 + 0.3*366 + 0.15*381

= 374.85

c) Using the exponential smoothing method, the forecast would be:

F(September 21) = 0.2*410 + 0.8*362

= 371.6

F(September 28) = 0.2*381 + 0.8*371.6

= 373.48

F(October 5) = 0.2*366 + 0.8*373.48

= 371.98

F(October 12) = 0.2*378 + 0.8*371.98

= 373.18